A tiny ball weighs 0.0055 kg and carries a charge of of 3.25 x10-6 C. What electric field (magnitude and direction) is needed for the ball to remain suspended in air?

 


1
Expert's answer
2022-02-18T08:22:41-0500

At the equilibrium (the ball remains suspended in air)

"mg=qE"

Hence, the electric field

"E=\frac{mg}{q}=\frac{0.0055*9.8}{3.25*10^{-6}}=1.66*10^4\:\rm V/m"

The field must be directed upward.